IPP Progress Report for the week YYYY.MM.DD - YYYY.MM.DD

(Up to IPP Progress Reports)

Eugene Magnier

  • finished restore of ipp023 - ipp026 data to their rightful location
  • ran 'delstar' on the big 3pi database to remove the (corrupted) SuperCOSMOS data
  • added Xoff,Yoff fields to the DVO schema to allow for measurement of the Koppenhofer effect
  • updated trunk with changes from my branch (mostly related to dvo improvements / schema issues)
  • worked on PS2 scenarios for hardware

Serge Chastel

  • Czar on Mon and Fri
  • Fixed ippMonitor/scripts/skycellplot.dvo: Having the same values in ptsize crashed the script (no image displayed by "Simple plot - staticsky").
  • More data for MOPS tests
  • Finished IPP-MOPS "ICD" (and waiting for feedback).

Heather Flewelling

  • worked on talk for ps1 seminar HeatherTalk
  • czar for tuesday and half of friday
  • worked on ipptopsps changes for sa11, see Sa10Problems?
  • worked on building sa11 (starting on friday, with new build)
  • czarred ThreePi?.V3 minidvodbs - there are a couple of problems to them due to the neb/other problems we experienced this week
  • worked on plan for LAP ingest into psps LAPintoPSPS

Mark Huber

  • MD04.refstack finally finished and SSdiffs caught up, MD05.refstack started
  • MD01.GR0 wrap-up processing, nightly stacks continue
  • czar typical issues, rebalanced stdscience/stack with stare, lm_sensor issue
  • PS seminar -- MDS+TSS

Bill Sweeney

  • spent most of the week working on lap static sky processing. The first batch of skycells triggered a few bugs in psphotStack. It's often slow going on those because the program runs so long. Fixed the problems. Finally, psphotStack is working well for skycells at low galactic latitude. After the bug fixes processing proceeded with only a couple of faults.
  • Debugged problem with the new stack detection id columns in psps object table. They are useless in sas 10 due to a bug in psps. Fixed that.
  • Enhanced the skycalResult table to include columns for the number of detections and the number of extended objects.
  • modified the staticsky script to run psphotStack on single filter runs
  • many night and weekend hours fixing problems with the cluster.

Chris Waters

  • Launched cleanup of old LAP stacks. This seems to have recovered ~100TB, largely on wave1 nodes.
  • MOPS:
    • Updated MOPS false positive tests wiki MOPS_diff_false_positives with results for unconvolved differences.
    • Identified that trail fitting issue is a result of the EXT_NSIGMA threshold for galaxies not working well for trailed sources.
    • Reprocessed data for Larry to compare processing results.
    • Sent camera stage magnitudes to Alan for comparison with diff results.
  • Background restoration
    • Planned and made database change to provide background_model column to manage background regeneration.
    • Wrote and began testing ippTask and ippScript to handle this regeneration.